Circular No. 6540

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVAE The Supernova Cosmology Project [cf. IAUC 6270, plus R. Knop and

P. Nugent at Cerro Tololo (CTIO); C. Lidman and M. Della Valle at the European Southern Observatory (ESO); and B. Schaefer with the WIYN telescope at Kitt Peak] report fourteen supernovae found with the CTIO 4-m telescope in their High Redshift Supernova Search. The last column indicates the discovery telescope: (1) Keck; (2) ESO 3.6-m reflector.

SN  1997 UT  α_{2000}  δ_{2000}  R  z  type G
1997F Jan. 5  4^{h}55^{m}14.^{s}27 − 5^{o}51′44′′.8 23.9 0.58 Ia (1)
1997G Jan. 5  4 58 30.23 − 3 16 04.0 23.7 0.75*a Ia (1)
1997H Jan. 5  4 59 36.57 − 3 09 34.6 22.8 0.53 Ia (2)
1997I Jan. 5  4 59 37.31 − 3 09 01.6 20.9 0.18* Ia (2)
1997J Jan. 5  7 41 17.83 + 9 33 11.4 23.4 0.62 Ia? (1)
1997K Jan. 6 7 54 55.09 + 4 19 01.3 23.6 0.59 ? (1)
1997L Jan. 5  8 21 57.14 + 3 53 35.7 23.1 0.55* Ia (2)
1997M Jan. 6 8 22 57.39 + 3 52 25.6 23.2 0.43 ? (1)
1997N Jan. 5  8 23 50.03 + 3 28 52.8 21.2 0.18* Ia (2)
1997O Jan. 6  8 24 02.51 + 4 07 33.1 23.7 0.37 Ia (1)
1997P Jan. 6 10 55 55.93 − 3 56 48.2 23.0 0.47 Ia (1)
1997Q Jan. 6 10 56 51.48 − 3 58 36.9 22.5 0.44 Ia (2)
1997R Jan. 6 10 57 19.23 − 3 54 51.7 24.4 0.65*a Ia (1)
1997S Jan. 6 10 57 51.59 − 3 45 46.9 23.6 0.61 Ia (1)

The supernovae were all discovered before or at maximum light, they were observed on Jan. 8–11 at CTIO and WIYN, and the supernova spectra were obtained at the Keck telescope on Jan. 12–13 and the ESO 3.6-m telescope on Jan. 13. The redshifts were determined from host-galaxy emission and absorption lines, except for those marked with an asterisk, which were determined solely from the spectra of the supernovae. In all cases with either a measured host-galaxy redshift or an adjacent-galaxy redshift (the latter indicated by ‘*a’), the supernova redshift is consistent. The supernovae are all within 2′′.5 of the host-galaxy center. Follow-up photometry is being made in R and I for the twelve supervovae at z > 0.35, and in V, R, and I for SNe 1997I and 1997N; finding charts are available from saul@lbl.gov.

1997 January 17  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green